OpenStudy (falling_in_katt):

***WILL MEDAL*** Solve this equation by completing the square. 3x^2 - 4x = -2

OpenStudy (chillout):

Do you know the steps to complete the square?

OpenStudy (falling_in_katt):

no...

OpenStudy (chillout):

All right. Here you go. Divide the whole equation by whatever is multiplying the squared term. In this cae you have to divide everything by 3. By doing this, you have \(x^{2}-\frac{4}{3}x=\frac{-2}{3}\)

OpenStudy (chillout):

Then you take half of the "x" coefficient (\(\frac{4}{3}\)), square it and add to both sides.

OpenStudy (chillout):

Which means you will be adding \((\frac{2}{3})^{2}\) to both sides.

OpenStudy (chillout):

By doing this, you will have \(x^{2}-\frac{4}{3}x+ \frac{4}{9}=\frac{-2}{3}+\frac{4}{9}\)

OpenStudy (chillout):

Do you see that the left side can be rewritten as a perfect square?

OpenStudy (falling_in_katt):

yes

OpenStudy (chillout):

And that's it! Rewrite the left side, simplify the right side and you will *ALMOST* have your answer. Can you finish it?

OpenStudy (falling_in_katt):

i can try

OpenStudy (chillout):

Take your chances :3

OpenStudy (falling_in_katt):

Im trying to square the left side and its not working

OpenStudy (chillout):

Well, rewriting it as a square we have \((x-\frac{2}{3})^{2}\)

OpenStudy (chillout):

check?

OpenStudy (falling_in_katt):

4/9?

OpenStudy (chillout):

Reminder that \((a-b)^{2}=a^{2}-2ab+b^{2}\) Moving on... We have \((x-\frac{2}{3})^{2}=\frac{-2}{9}\) and \(x-\frac{2}{3}=±\sqrt{\frac{-2}{9}}\) This problem WILL NOT have answers in the set of reals, ONLY in the set of complex numbers.

OpenStudy (falling_in_katt):

Oh. That makes more sense.

OpenStudy (chillout):

Does it ask for its roots?

OpenStudy (falling_in_katt):

No

OpenStudy (chillout):

Well, then there you have your answer. :D

OpenStudy (falling_in_katt):

Thank you